OpenStudy (anonymous):

Amy paid $76.22 for a pair of shoes during a 40% off sale. What was the reg. price?

OpenStudy (amistre64):

still too much ..

OpenStudy (amistre64):

76.22 = P - P(.4) 76.22 = P(1-.4) 76.22 ----- = P .6

OpenStudy (anonymous):

amistre has it, but i love to butt in. if you got 40% off it means you paid 60% of the price. \[76.22=.6P\] \[P=\frac{76.22}{.6}\]

OpenStudy (amistre64):

lol..... same diff